Water is flowing at a rate of 100 L/s in a 20 cm diameter pipe. How fast is the water flowing in m/s?


1
Expert's answer
2021-11-22T10:11:23-0500

"Q=vS,\implies"

"v=\frac QS=\frac{4Q}{\pi d^2}=3.2~\frac ms."